OpenStudy (anonymous):

Need help solving 0 = −x^2 + 2.5x + 1 I have this formula, but I have no idea how to plug it in or how to write it down or what to do with the 2.5 Any help would be MUCH appreciated. H(t) = -16^2 + vt + s

OpenStudy (anonymous):

\[ 0 = −x^2 + 2.5x + 1\] \[10x^2-25x+10=0\] \[2x^2-5x+2=0\]now if you can factor, then factor, if not use the quadratic formula

OpenStudy (anonymous):

this one actually factors as \((x-2)(2x-1)=0\) so the rest should be ok

OpenStudy (anonymous):

I don't understand how you got 10x^2, 25, and 10. Did you multiply by 10?

Directrix (directrix):

@satellite multiplied every member of the first equation by 10 to clear the decimals.

OpenStudy (anonymous):

i hate decimals

OpenStudy (anonymous):

then divide by the common factor of 5

OpenStudy (anonymous):

When you factor, aren't you supposed to find the product of the last term and the sum of the middle term? When I did that, it said that -2 *-1 = 2 but 2 + -1 = -3 I checked it on Wolfram and it said that (x−2)(2x−1) was right, but is there something I'm missing?

OpenStudy (anonymous):

yes, the leading coefficient is 2, not 1, so that whole "sum product" business needs to be readjusted

OpenStudy (anonymous):

I've been trying to write and rewrite the equation from factoring but even when I rearrange it, I can't get -5 from -1 and -2.